### v2

# Beyond Attraction: A Novel Approach to Repulsive Casimir-Lifshitz Forces using Controllable Off-stoichiometry in Gapped Metals

Abstract
We uncover a novel physical mechanism that enables a switch between attractive and repulsive Casimir forces when a Teflon surface interacts with a new form of quantum material (i.e., gapped metal) surface across different liquid media. We demonstrate the discovery of a zero-frequency Casimir effect, which, for the first time, reveals the potential for quantum switching within nanometer distances-a scale previously thought to be unattainable. Hence, our results introduce a new method to induce phase (stoichiometry)-controlled attraction-repulsion transitions and achieving quantum levitation in a liquid medium by tuning the liquid environment. This study thus not only advances our understanding of quantum forces at the nanoscale via their correlation to dielectric properties of involved materials but also opens up exciting possibilities for their manipulation in novel ways, forming the basis towards innovative advancements in nanoscale technology.
I Introduction
The potential to have repulsive Casimir-Lifshitz forces was implicit already in the original Lifshitz theory Dzyaloshinskii et al. (1961) and conclusively demonstrated in the 1970s Anderson and Sabisky (1970); Hauxwell and Ottewill (1970); Richmond and Ninham (1971); Richmond et al. (1973). Past investigations demonstrating repulsive Casimir-Lifshitz forces have two important features: first, retardation effects are expected at relatively large distances (3-100 nm and beyond) Boström et al. (2012); Esteso et al. (2015, 2016); Zhao et al. (2019); Esteso et al. (2022), and second, the zero frequency contribution to the Casimir force is only significant at much, much longer distances ($\gtrsim$ 100 nm) Boström and Sernelius (2000).
In this work, we propose a novel approach based on recent developments in understanding the properties of gapped metalsMalyi and Zunger (2020); Khan et al. (2023). These materials have a Fermi level located inside the conduction (or less often valence) bands, resulting in a high free carrier concentration in compounds with large internal band gaps between the principal band edges. Due to such unique electronic structure, they could exhibit a spontaneous off-stoichiometry. For instance, in the case of n-type gapped metals, the formation of acceptor cation vacancy can result in the decay of conducting electrons to the defect state, resulting in the lowering defect formation energy Malyi et al. (2019). Due to such unique physical phenomena, the properties of gapped metals are significantly distinct between different stoichiometries, often with a range of stable compounds that can be synthesized by tuning environmental conditions. Using the unique properties of these materials, we propose a novel approach that enables us to control the sign of the Casimir-Lifshitz free energy, by moving between surface regions with different optical properties, to avoid stiction problems, equivalent to a simple method to turn a nanoscale switch on and off, based on off-stoichiometric properties of a gapped metal substrate interacting with a teflon nanosurface in fluids. This leads to observable sign changes, as well as retardation and entropic effects, at separations as small as a few nanometers.
Figure 1: (Colors online) (a) Schematic figure of the three-layer system involved here: gapped metal of infinite thickness and dielectric function $\varepsilon_{1}$, in contact with a liquid layer of dielectric function $\varepsilon_{2}$, of thickness $d$, which is in turn in contact with PTFE of dielectric function $\varepsilon_{3}$. (b) The dielectric functions at imaginary frequencies $\xi_{m}$ for different Ca6-xAl7O16 compounds (note that not all compounds are gapped metals as at high Ca vacancy concentration they become insulators), liquids (Iodobenzene, Bromobenzene, Chlorobenzene and Methanol), and PTFE surface. Zero-frequency dielectric functions of all these materials and liquids are as follows, $\varepsilon_{\mathrm{Ca_{6}Al_{7}O_{16}}}(0)$ = 300.4216, $\varepsilon_{\mathrm{Ca_{5.75}Al_{7}O_{16}}}(0)$ = 58.4389, $\varepsilon_{\mathrm{Ca_{5.5}Al_{7}O_{16}}}(0)$ = 2.815, $\varepsilon_{\mathrm{Iodobenzene}}(0)$ = 4.6, $\varepsilon_{\mathrm{Bromobenzene}}(0)$ = 5.37, $\varepsilon_{\mathrm{Chlorobenzene}}(0)$ = 5.75, $\varepsilon_{\mathrm{Methanol}}(0)$ = 32.9, $\varepsilon_{\mathrm{PTFE}}(0)$ = 2.1 .
Notably, a method is shown where “switching" can be achieved at a fixed distance by varying between different surface patches with different optical properties. We demonstrate that this can be achieved both at the extreme nanoscales where additional non-Casimir-Lifshitz forces originating from the fluid are expected to impact the measured force when the surfaces are a few nanometers apart and at the submicron scales. The retarded Casimir-Lifshitz interaction depends on an intricate balance between repulsive contributions from low frequencies ($\lesssim 10^{15}$ rad/s) and attractive contributions from higher frequency ranges, leading to the surprising importance of zero-frequency contributions at small distances, contrary to what is reported in the literature. Moreover, correlations of dielectric functions between different phases of gapped metals in the fully retarded theory influence a potential attractive-repulsive transition at 2 nm and 5.9 nm for metallic phases of gapped metals, respectively. Past work described the repulsion between gapped metal surfaces and air bubbles Boström et al. (2023) and attraction between a heterogeneous gapped metal surface and a gold surface Boström et al. (2024). In the current work, we will explore a novel method for tuning attraction/repulsion transitions via combined retardation and entropic effects in previously unattainable nanolimits. One of the key elements to achieve repulsive Casimir interaction is the choice of intervening medium (i.e. liquid) between two interacting materials. In the case of repulsive Casimir interaction, the necessary conditions $\varepsilon_{1}>\varepsilon_{2}>\varepsilon_{3}$, where $\varepsilon_{1}$, $\varepsilon_{3}$, and $\varepsilon_{2}$ represents the dielectric response of the two materials and the intervening liquid, respectively, must be satisfied. As we will demonstrate, exploiting the properties of gapped metals makes it possible to have both separation and phase change-induced attraction/repulsion transitions. Furthermore, from a fundamental point of view, it is of interest that retardation, the closeness of dielectric functions between liquids and teflon, and zero frequency contribution are connected enabling us to predict how to tune off-stoichiometrically controlled attraction/repulsion switches at unprecedented small separation ranges. To our knowledge, the striking dominance of the zero-frequency contribution at nanoscale separations and low temperatures has not been previously reported. Notably, this can be manipulated by varying the separation between metal and PTFE and interacting regions having different optical properties.
II Theory
We consider the three-layer planar system reported in Fig. (1a) as our model system , where we calculate the Casimir-Lifshitz interaction free energy $F(d)$ between a planar gapped metal (medium 1), dielectric functions $\varepsilon_{1}(i\xi_{m})$, and a planar surface (medium 3) at a specific temperature $T=300$ K from the calculated dielectric functions at discrete imaginary frequencies $\xi_{m}$ in Fig. (1b). The medium 3 is considered to be made of Teflon (PTFE) van Zwol and Palasantzas (2010). The intervening medium 2 is of the different liquids (Methanol, Bromobenzene, Iodobenzene, and Chlorobenzene) with dielectric functions $\varepsilon_{2}(i\xi_{m})$. The dielectric functions of Teflon and the different liquids are taken from van Zwol and Palasantzas van Zwol and Palasantzas (2010). The dielectric functions of the gapped metals were obtained from Table (1). Lebedew was the first to realize the connection between intermolecular forces and radiation processes in 1894 Lebedew (1894). The theory confirming this relation of optics with forces was presented in the classic paper by Lifshitz and co-workers Dzyaloshinskii et al. (1961). The Casimir-Lifshitz interaction free energy can be written in terms of the spectral function, $\mathrm{g}(\xi_{m})$, as Ninham et al. (1970); Ninham and Lo Nostro (2010); Sernelius (2018),
$F(d)$ $={\sum_{m=0}^{\infty}}{}^{\prime}\mathrm{g}(\xi_{m})$ (1)
[…]
with $F(d)=-A(d)/(12\pi d^{2})$ and where $\sigma=\rm TE,TM$, and the prime in the sum above indicates that the first term ($m$ = 0) has to be weighted by $1/2$. The Fresnel reflection coefficients between surfaces $i$ and $j$ for the transverse magnetic (TM) and transverse electric (TE) polarizations are given by
$r_{\rm TE}^{ij}=\frac{\kappa_{i}-\kappa_{j}}{\kappa_{i}+\kappa_{j}}\,;\,\,\,\,\,r_{\rm TM}^{ij}=\frac{\varepsilon_{j}\kappa_{i}-\varepsilon_{i}\kappa_{j}}{\varepsilon_{j}\kappa_{i}+\varepsilon_{i}\kappa_{j}}\,.$ (2)
Here $\kappa_{i}=\sqrt{{q}^{2}+\varepsilon_{i}\xi_{m}^{2}/c^{2}}$, with $i=1,2,3$ and the Matsubara frequency being $\xi_{m}=2\pi kTm/\hbar$, where $c$ and $k_{B}$ are the velocity of light in vacuum and the Boltzmann constant. For our system, a balance between repulsive contributions in the low-frequency range with high-frequency attractive contributions leads to a possible sign change in the Casimir-Lifshitz interaction when comparing the non-retarded results with the fully retarded theory. The origin of this effect is due to the distance-dependent exponential of the spectral function $\mathrm{g}(\xi_{m})$ in Eq. (1). For fluid systems, additional short-range structural potentials can exist at interfaces. For liquids with isotropic or anisotropic molecules, the granularity near interfaces can give rise to attractive or repulsive oscillatory forces. For up to at least 6 molecular layers, these forces can invalidate the use of the continuum Lifshitz theory Qu et al. (2007); Bonaccurso et al. (2008); Ninham and Lo Nostro (2010). However, surface roughness can do two things: it can induce uncertainty in the distance and, potentially, smoothen out oscillations. Short-range forces were recently explored for water/ice systems. These give rise to repulsive forces acting over distances of the order of at least a few molecular diameters. These interactions originate from corrections to the bulk density profile near a solid-fluid interface. In the case of ice-water-vapor, it could be estimated in the following way Luengo-Marquez et al. (2022),
$F_{sr}(d)=B_{2}e^{-\kappa_{2}d}-B_{1}e^{-\kappa_{1}d}cos(q_{0}d+\phi)$ (3)
The parameters can be derived using molecular simulations Luengo-Marquez et al. (2022). We will in the following ignore similar contributions for our liquid systems. The focus is on exploring the previously measured attractive or repulsive contributions from Casimir-Lifshitz interactions Lee and Sigmund (2002); Feiler et al. (2008); Munday et al. (2009); Zhao et al. (2019) at the nano to submicron scale tuned by the optical properties of gapped metals.
III Results
We consider here a three-layer Ca6-xAl7O16–Methanol–PTFE interface. The metallic nature of gapped material is observed in Ca6Al7O16 and Ca5.75Al7O16, while Ca5.5Al7O16 represents the insulating phase of gapped material. To explain our results, we first focus on the short-range limit. In the non-retarded limit ($c\to\infty$), the Fresnel reflection coefficients for transverse electric (TE) polarization approach zero, and the remaining contribution comes entirely from the transverse magnetic (TM) light polarization,
$\mathrm{r}_{\rm NR}^{12}\mathrm{r}_{\rm NR}^{32}=\frac{\varepsilon_{1}-\varepsilon_{2}}{\varepsilon_{1}+\varepsilon_{2}}\times\frac{\varepsilon_{3}-\varepsilon_{2}}{\varepsilon_{3}+\varepsilon_{2}}\,.$ (4)
Hence, the Hamaker constant (in the non-retarded limit) can be expressed as,
$A^{\text{NR}}=\frac{3}{2}k_{B}T{\sum_{m=0}^{\infty}}{}^{\prime}\sum_{j=1}^{\infty}\Big{[}r_{\text{NR}}^{21}(i\xi_{m})r_{\text{NR}}^{23}(i\xi_{m})\Big{]}^{j}/j^{3}.$ (5)
There are crossovers between the product of the reflection coefficients in the non-retarded limit for different stoichiometry of the gapped Ca6-xAl7O16 metal surfaces in Fig. (2), that drives the possibility of attraction to repulsion transition in Casimir-Lifshitz interaction.
Figure 2: (Colors online) Product of the reflection coefficients at imaginary frequencies $\xi_{m}$ in non-retarded limit for ranges of gapped metals, Ca6-xAl7O16.
Hence, in a broad frequency range $\mathrm{r}_{\rm NR}^{12}\mathrm{r}_{\rm NR}^{32}>0$, the non-retarded Hamaker constant in Eq. (5) is positive, corresponding to an attractive interaction. However, other frequency intervals produce repulsive contributions.
Figure 3: (Colors online) The spectral functions $\mathrm{g}(\xi_{m})$, defined in Eq. (1), at different separations of liquid layer thickness in comparing the fully retarded and non-retarded limits in Ca6-xAl7O16-Methanol-PTFE like systems. Solid lines indicate the retardation effect, while dashed lines represent the behavior of spectral functions in the non-retarded limit.
Figure 4: (Colors online) (a)Free energy vs. distance plot in Ca6-xAl7O16 – Methanol – PTFE like system. Ca6-xAl7O16 indicates different stoichiometry. Here red color dot and magenta color box denote maximum value of Casimir energy for Ca6Al7O16 and Ca5.75Al7O16 are occurring at 2.9 nm and 9.49 nm respectively. (b) Decomposition of the contributions to the total non-retarded (non-ret) and retarded (ret) Casimir-Lifshitz free energy for Ca6Al7O16–Methanol–PTFE like system.
In Fig. (3), we present the decomposition of the Casimir-Lifshitz free energy into distance and frequency-dependent spectral functions, $\mathrm{g}(\xi_{m})$. At large separations of liquid layer thickness, retardation effect influences the contributions from all frequencies except for the extremely low ones. Ultimately, as it is well known Dzyaloshinskii et al. (1961); Ninham et al. (1970); Ninham and Daicic (1998); Boström and Sernelius (2000), only the zero-frequency term survives. As the separation decreases, higher and higher frequencies play a role. As a result, the effect of retardation on repulsive contributions for the smallest separations is smaller since these originate from the low-frequency range. The key idea of this effect is a product of reflection coefficients (discussed in the previous part) along with a distance-dependent exponential factor $e^{-2\sqrt{{q}^{2}+\varepsilon_{2}\xi_{m}^{2}/c^{2}}d}$, meaning that significant contributions appear only when $\xi_{m}\sim\mathcal{O}(1/d)$ or smaller.
In Fig. (4), we notice that the non-retarded Casimir-Lifshitz interaction for different stoichiometry of Ca6-xAl7O16 is attractive in all separations, but there are reversal of signs for the retarded interaction around 2 nm and 5.9 nm for the metallic phases of the gapped metals Ca6Al7O16 and Ca5.75Al7O16, respectively. This sign reversal can also be viewed from zero-frequency Hamaker constant in Table. (2). In order to demonstrate the long-range characteristics of various phases of gapped metals, the energy variation in the limit of long distances is illustrated in Fig. (5). It is observed that the metallic phases of gapped metals exhibit a repulsive effect even at long distances while the insulating phase displays an attractive interaction.
modes ($\omega_{j}$) Coefficients ($C_{j}$) for different Ca6-xAl7O16 compounds
Ca6Al7O16 Ca5.75Al7O16 Ca5.5Al7O16
0.0206 58.9601 0.6494 0.0001
0.0347 91.1774 1.797 0.0003
0.0587 57.4068 5.2997 0.001
0.1013 16.4729 15.4951 0.0221
0.1996 73.0451 29.8463 0.3283
0.3938 0.3949 2.1519 0.7511
0.9556 0.0706 0.2519 0.4345
2.2773 0.0987 0.0392 0.0
6.4732 0.4594 0.2384 0.2279
10.2048 0.7938 0.8189 0.0
18.2421 0.3705 0.474 0.0486
30.9018 0.1655 0.2388 0.0
54.455 0.0059 0.0283 0.001
Table 1: Parametrization of the average dielectric function of continuous media, $\varepsilon(i\xi)$, at imaginary frequencies for Ca6-xAl7O16 as calculated with first-principles calculations and a damping coefficient ($\Gamma$) set to 0.2  eV. The $\omega_{j}$ modes are given in $\rm eV$. The largest difference between fitted and calculated $\varepsilon(i\xi)$ is 0.08%.
Material $A^{NR}$ $A_{m=0}$
Ca6Al7O16 0.004 eV -0.012 eV
Ca5.75Al7O16 0.011 eV -0.004 eV
Ca5.5Al7O16 0.033 eV 0.016 eV
Table 2: The non-retarded Hamaker constant for gapped metals–methanol–PTFE combinations.
Figure 5: (Colors online) Free energy vs.distance plot in Ca6-xAl7O16 – Methanol – Teflon like system in the long-distance limit. Ca6-xAl7O16 indicates different stoichiometry
The dielectric functions are larger for the metallic phases than for the insulating phase in the low-frequency range leading to the enhancement of repulsive contributions. In this figure, there is a smooth transition between the attractive van der Waals regime and the long-range Casimir region around 2 nm and 5.9 nm for the metallic phases of the gapped metals.
Remarkably, in most metallic cases the zero-frequency term dominates at 2.9 nm due to retardation and a strong cancellation of attractive and repulsive contributions. To the best of our knowledge, zero-frequency effects have never been predicted before, or observed, at such small distances at room temperature. We observe that the long-range Casimir asymptote results in a repulsive effect originating from the presence of zero-frequency transverse magnetic (TM) modes and the influence of retardation. These results are consistent with Fig. (3). The contribution from TM modes is dominant for all distances but there is a small impact of TE modes, especially around 1.5 nm  where TM modes are canceling out in Fig. (4b). If we discard the impact of retardation, the contributions from these modes are absent, as shown in Fig. (4). We see exactly the opposite trends in the non-retarded limit in Fig. (6) when the intervening medium is considered to be Iodobenzene instead of Methanol. The dielectric functions for Iodobenzene, PTFE and insulating phase of gapped metal, Ca5.5Al7O16 lead to a sign reversal around 75 nm in Casimir energy as the distance varies and the observed long-range attraction. The explanation is very straightforward: the repulsive interaction is linked to the short-range Hamaker constant crossing over to attraction when retardation effect mitigates the influence of finite frequency terms. The attraction is connected to the long-range zero frequency Hamaker constant. In contrast, for the other stoichiometries, the interaction remains repulsive across all distances. As shown in the inset of Fig. (6), a mainly zero-frequency driven energy minima is predicted around 0.11 $\mu$m. This leads to the possibility of trapping a PTFE nanoparticle in iodobenzene near the surface of the gapped metal in the insulating phase. Similar analysis was also carried out using various liquids such as Bromobenzene and Chlorobenzene. Analogous patterns to those observed with Iodobenzene are identified, suggesting potential trapping of PTFE nanoparticles at the interface of the insulating layer of the gapped materials.
Figure 6: (Colors online) Free energy vs. distance plot in Ca6-xAl7O16 –Iodobenzene– PTFE like system. (a) In the inset, we illustrate the long-range characteristics of the Casimir-Lifshitz free energy, with an emphasis on the predominant contribution from zero-frequency term.
IV Conclusions
Here, we have reported the development of a theoretical model that predicts how moving across regions with different off-stoichiometry can be used to tailor the sign of the Casimir-Lifshitz force between the Teflon nanosurface and gapped metal surfaces separated by methanol. Specifically, considering the range of gapped metals Ca6-xAl7O16, we demonstrate that controllable off-stoichiometry regions can be used to tune the Lifshitz interaction from attraction to repulsion. This behavior originates from the effect of off-stoichiometry on the dielectric function of gapped metals - the formation of cation vacancies induces a reduction of free carriers and the Drude contribution to the dielectric function. We show that the phase transition from insulator (Ca5.5Al7O16) to metal (Ca6Al7O16 and Ca5.75Al7O16) leads to a change in the force from attraction to repulsion at the nano to sub-micron scale driven by retardation and caused by the difference in dielectric functions of the gapped metals. The key point in our work is that a zero-frequency contribution and its interconnection with retardation and closeness of dielectric function of the liquids media and PTFE are the main origin for the sign change switching at unprecedented small distances. To the best of our knowledge, in the past purely zero-frequency effects have always been assumed to occur at much larger separations or at much higher temperatures Ninham and Daicic (1998). Here, we see its domination for separations as small as a few nanometers in the metallic case but not for the insulating phase. While the dielectric functions used here do not have any temperature dependence (and hence can’t be used to represent the real temperature-dependent picture), one can find that, e.g., 10% changes in temperature, and the corresponding changes in the zero-frequency term, could lead to a measurable change in attraction-repulsion transition distances. With a change in background liquid media, it is also possible to have a zero-frequency Casimir-Lifshitz term leading to quantum levitation. In this way, using different liquids, we propose a new tool for nanoengineering via phase-transition-related quantum levitation in a liquid medium.
Acknowledgements.
SP, OIM, and MB’s research contributions are part of the project No. 2022/47/P/ST3/01236 co-funded by the National Science Centre and the European Union’s Horizon 2020 research and innovation programme under the Marie Skłodowska-Curie grant agreement No. 945339. The research by SP and MB took place at the “ENSEMBLE3 - Centre of Excellence for nanophotonics, advanced materials and novel crystal growth-based technologies" project (grant agreement No. MAB/2020/14) carried out within the International Research Agendas programme of the Foundation for Polish Science co-financed by the European Union under the European Regional Development Fund, the European Union’s Horizon 2020 research and innovation programme Teaming for Excellence (grant agreement. No. 857543) for support of this work. S.O. thanks the National Science Centre, Poland (grant no. UMO/2020/39/I/ST4/01446) and the “Excellence Initiative - Research University" (IDUB) Program, Action I.3.3 - “Establishment of the Institute for Advanced Studies (IAS)" for funding (grant no. UW/IDUB/2020/25). We gratefully acknowledge Poland’s high-performance computing infrastructure PLGrid (HPC Centers: ACK Cyfronet AGH) for providing computer facilities and support within computational grant no. PLG/2023/016228 and for awarding this project access to the LUMI supercomputer, owned by the EuroHPC Joint Undertaking, hosted by CSC (Finland) and the LUMI consortium through grant no. PLL/2023/4/016319. M.B. gratefully acknowledges the kind hospitality at the University of Cagliari where this work was completed.