### v3

# Beyond Attraction: A Novel Approach to Repulsive Casimir-Lifshitz Forces using Heterogeneous Off-stoichiometry in Gapped Metals

Abstract
We uncover a physical system that enables a switch between attractive and repulsive Casimir forces when a Teflon surface interacts with a new form of quantum material (i.e., gapped metal) surface across different liquid media. Past works usually suggest that entropic effects on the Casimir force (originating partly from its zero frequency contribution that has an approximately linear temperature dependence) occur at high temperatures and/or large separations. We propose a new way to achieve a zero-frequency Casimir effect, which reveals the potential for a Casimir force quantum switching within nanometer distances-a scale often thought to be unattainable. Furthermore, the heterogeneous surface can have different phases (both metallic and insulating). Hence, our results introduce a new method to induce phase (stoichiometry)-controlled attraction-repulsion transitions and to achieve quantum levitation in a liquid medium by tuning the liquid environment. This study opens up new possibilities for manipulating the Casimir interactions between surfaces, forming the basis for advancements in nanoscale technology.
keywords:
Casimir Physics; Attraction-Repulsion Transition; Entropic Effects; Heterogeneous quantum materials
journal: Physics Letters A
[…]
[label1]organization=Universita degli Studi di Palermo, Dipartimento di Fisica e Chimica–Emilio Segre, addressline=Via Archirafi 36, city=Palermo,postcode=90123, country=Italy
[…]
[label2]organization=Centre of Excellence ENSEMBLE3, addressline=Wolczynska Str. 133, city=Warsaw, postcode=01-919, country=Poland
[…]
[label3]organization=Chemical and Biological Systems Simulation Lab, Centre of New Technologies, University of Warsaw, addressline=Banacha 2C, city=Warsaw, postcode=02-097, country=Poland
[…]
[label5]organization=Qingyuan Innovation Laboratory,city=Quanzhou,postcode=362801,country=China
{graphicalabstract}[Uncaptioned image]
{highlights}
Casimir effect for heterogeneous phase change materials
Entropy effects on Casimir forces at nanometer scales
Proposed Casimir switch in fluids for nanotechnological applications
1 Introduction
The potential to have repulsive Casimir-Lifshitz forces was implicit already in the original Lifshitz theory [1] and conclusively demonstrated in the 1970s [2, 3, 4, 5]. Past investigations demonstrating repulsive Casimir-Lifshitz forces have two important features: first, retardation effects are expected at relatively large distances (3-100 nm and beyond) [6, 7, 8, 9, 10], and second, the zero frequency contribution to the Casimir force is usually only significant at much, much longer distances ($\gtrsim$ 100 nm) [11, 12, 13, 14, 15, 16]. The engineered switching of Casimir force between attraction to repulsion has in the past been reported using topological insulators [17, 18], tin [19], and Vanadium dioxide [20, 21]. It is also worth mentioning that in a very recent paper [22] , the Casimir force contribution from the zero frequency term gave, similar to our results, an important contribution at a short separations smaller than 100 nm. This is because the the closeness of dielectric functions between solution and the coated nanofilms at high frequency, while the contrast of dielectric permittivities is high at zero frequency. As a positive derivative of the free energy per unit area corresponds to repulsive pressure, the system in some, but not all, cases considered in the current work is unstable, i.e., the gap distance is a little larger than the zero-force position, and the top and bottom layers get far away, which is different from stable systems prepared by Zhao et al.[9] Novel control of the switching from attractive to the repulsive forces by twist has in the past included exploiting layers of lithium iodate (LiIO3), separated by a composite medium of nanoparticles dispersed in a liquid [23] A detailed review of the current state of Casimir effects and its versatile applications has been given by Klimchitskaya et al.[24].
In this work, we propose a novel approach based on recent developments in understanding the properties of gapped metals[25, 26]. These materials have a Fermi level located inside the conduction (or less often valence) bands, resulting in a high free carrier concentration in compounds with large internal band gaps between the principal band edges. Due to such a unique electronic structure, they could exhibit spontaneous off-stoichiometry. For instance, in the case of n-type gapped metals, the formation of acceptor cation vacancy can result in the decay of conducting electrons to the defect state, lowering defect formation energy [27]. Due to such unique physical phenomena, the properties of gapped metals are significantly distinct between different stoichiometries, often with a range of stable compounds that can be synthesized by tuning environmental conditions. Using the unique properties of these materials, we propose a novel approach that enables us to control the sign of the Casimir-Lifshitz free energy, by moving between surface regions with different optical properties to avoid stiction problems, equivalent to a simple method to turn a nanoscale switch on and off, based on off-stoichiometric properties of a gapped metal substrate interacting with a teflon nanosurface in fluids. This leads to observable sign changes, as well as retardation and entropic effects, at separations as small as a few nanometers.
Figure 1: (Colors online) Schematic figure of the three-layer system involved here: gapped metal of infinite thickness and dielectric function $\varepsilon_{1}$, in contact with a liquid layer of dielectric function $\varepsilon_{2}$, of thickness $d$, which is in turn in contact with PTFE of dielectric function $\varepsilon_{3}$.
Figure 2: (Colors online) The dielectric functions at imaginary frequencies $\xi_{m}$ for different Ca6-xAl7O16 compounds (note that not all compounds are gapped metals as at high Ca vacancy concentration they become insulators), liquids (Iodobenzene, Bromobenzene, Chlorobenzene and Methanol), and PTFE surface. Zero-frequency dielectric functions of all these materials and liquids are as follows, $\varepsilon_{\mathrm{Ca_{6}Al_{7}O_{16}}}(0)$ = 300.4216, $\varepsilon_{\mathrm{Ca_{5.75}Al_{7}O_{16}}}(0)$ = 58.4389, $\varepsilon_{\mathrm{Ca_{5.5}Al_{7}O_{16}}}(0)$ = 2.815, $\varepsilon_{\mathrm{Iodobenzene}}(0)$ = 4.6, $\varepsilon_{\mathrm{Bromobenzene}}(0)$ = 5.37, $\varepsilon_{\mathrm{Chlorobenzene}}(0)$ = 5.75, $\varepsilon_{\mathrm{Methanol}}(0)$ = 32.9, $\varepsilon_{\mathrm{PTFE}}(0)$ = 2.1 .
Notably, a method is shown where “switching” can be achieved at a fixed distance by varying between different surface patches with different optical properties. We demonstrate that this can be achieved both at the extreme nanoscales where additional non-Casimir-Lifshitz forces originating from the fluid are expected to impact the measured force when the surfaces are a few nanometers apart and at the submicron scales. The retarded Casimir-Lifshitz interaction depends on an intricate balance between repulsive contributions from low frequencies ($\lesssim 10^{15}$ rad/s) and attractive contributions from higher frequency ranges, leading to the surprising importance of zero-frequency contributions at small distances, contrary to what is reported in the literature. Moreover, correlations of dielectric functions between different phases of gapped metals in the fully retarded theory influence a potential attractive-repulsive transition at 2 and 5.9 nm for metallic phases of gapped metals, respectively. Past work described the repulsion between gapped metal surfaces and air bubbles [28] and attraction between a heterogeneous gapped metal surface and a gold surface [29]. Herein, we explore a novel method for tuning attraction/repulsion transitions via combined retardation and entropic effects in previously unattainable nanolimits. One of the key elements to achieve repulsive Casimir interaction is the choice of intervening medium (i.e. liquid) between two interacting materials. In the case of repulsive Casimir interaction, the necessary conditions $\varepsilon_{1}>\varepsilon_{2}>\varepsilon_{3}$, where $\varepsilon_{1}$, $\varepsilon_{3}$, and $\varepsilon_{2}$ represents the dielectric response of the two materials and the intervening liquid, respectively, must be satisfied. As we will demonstrate, exploiting the properties of gapped metals makes it possible to have both separation and phase change-induced attraction/repulsion transitions. Furthermore, from a fundamental point of view, it is of interest that retardation, the closeness of dielectric functions between liquids and Teflon, and zero frequency contribution are connected, enabling us to predict how to tune off-stoichiometrically controlled attraction/repulsion switches at unprecedented small separation ranges. To our knowledge, the striking dominance of the zero-frequency contribution at nanoscale separations and low temperatures has not been previously reported. Notably, this can be manipulated by varying the separation between metal and PTFE and interacting regions having different optical properties.
2 Theory
We consider the three-layer planar system reported in Fig. (1) as our model system, where we calculate the Casimir-Lifshitz interaction free energy $F(d)$ between a planar gapped metal (medium 1), dielectric functions $\varepsilon_{1}(i\xi_{m})$, and a planar surface (medium 3) at a specific temperature $T=300$ K from the calculated dielectric functions at discrete imaginary frequencies $\xi_{m}$ in Fig. (2). The medium 3 is considered to be made of Teflon ( Polytetrafluoroethylene, PTFE) [30]. The intervening medium 2 is of the different liquids (Methanol, Bromobenzene, Iodobenzene, and Chlorobenzene) with dielectric functions $\varepsilon_{2}(i\xi_{m})$. The dielectric functions of Teflon and the different liquids are taken from van Zwol and Palasantzas [30]. We use DFT-calculated dielectric functions for the experimentally available Ca6-xAl7O16 compounds [28, 29]. We develop a parametrization of the average dielectric function for each of the considered compounds using an oscillator model:
$\varepsilon(i\xi)=1+\sum_{j}\frac{C_{j}}{1+(\xi/\omega_{j})^{2}}.$ (1)
Here, $\omega_{j}$ and $C_{j}$ represent the characteristic frequencies and the oscillator strength, respectively. The parameters can be obtained from Table (1). Lebedew was the first to realize the connection between intermolecular forces and radiation processes in 1894 [31]. The theory confirming this relation of optics with forces was presented in the classic paper by Lifshitz and co-workers [1]. The Casimir-Lifshitz interaction free energy can be written in terms of the spectral function, $\mathrm{g}(\xi_{m})$, as [32, 33, 34],
$F(d)$ $={\sum_{m=0}^{\infty}}{}^{\prime}\mathrm{g}(\xi_{m})$ (2)
$={\sum_{m=0}^{\infty}}{}^{\prime}\ \frac{k_{B}T}{2\pi}\int\limits_{0}^{\infty}dq\,q\sum_{\sigma}\ln(1-r_{\sigma}^{21}r_{\sigma}^{23}\mathrm{e}^{-2\kappa_{2}d})$
with $F(d)=-A(d)/(12\pi d^{2})$ and where $\sigma=\rm TE,TM$, and the prime in the sum above indicates that the first term ($m$ = 0) has to be weighted by $1/2$. The Fresnel reflection coefficients between surfaces $i$ and $j$ for the transverse magnetic (TM) and transverse electric (TE) polarizations are given by
$r_{\rm TE}^{ij}=\frac{\kappa_{i}-\kappa_{j}}{\kappa_{i}+\kappa_{j}}\,;\,\,\,\,\,r_{\rm TM}^{ij}=\frac{\varepsilon_{j}\kappa_{i}-\varepsilon_{i}\kappa_{j}}{\varepsilon_{j}\kappa_{i}+\varepsilon_{i}\kappa_{j}}\,.$ (3)
Here $\kappa_{i}=\sqrt{{q}^{2}+\varepsilon_{i}\xi_{m}^{2}/c^{2}}$, with $i=1,2,3$ and the Matsubara frequency being $\xi_{m}=2\pi kTm/\hbar$, where $c$ and $k_{B}$ are the velocity of light in vacuum and the Boltzmann constant. For our system, a balance between repulsive contributions in the low-frequency range with high-frequency attractive contributions leads to a possible sign change in the Casimir-Lifshitz interaction when comparing the non-retarded results with the fully retarded theory. The origin of this effect is due to the distance-dependent exponential of the spectral function $\mathrm{g}(\xi_{m})$ in Eq. (2). For fluid systems, additional short-range structural potentials can exist at interfaces. For liquids with isotropic or anisotropic molecules, the granularity near interfaces can give rise to attractive or repulsive oscillatory forces. For up to at least 6 molecular layers, these forces can invalidate the use of the continuum Lifshitz theory [35, 36, 37]. However, surface roughness can do two things: it can induce uncertainty in the distance and, potentially, smoothen out oscillations. Here it is important to recall that surface roughness can impact both the physics of nanoresonators [38] and the measured forces [39]. However, we will ignore structural and roughness contributions in our calculations. The focus is rather on exploring effects on the nanoscale that have previously been measured for many orders of magnitude larger separations [9, 40, 41, 42].
3 Results
We consider here a three-layer Ca6-xAl7O16–Methanol–PTFE interface. The metallic nature of gapped material is observed in Ca6Al7O16 and Ca5.75Al7O16, while Ca5.5Al7O16 represents the insulating phase of gapped material. To explain our results, we first focus on the short-range limit. In the non-retarded limit ($c\to\infty$), the Fresnel reflection coefficients for transverse electric (TE) polarization approach zero, and the remaining contribution comes entirely from the transverse magnetic (TM) light polarization,
$\mathrm{r}_{\rm NR}^{12}\mathrm{r}_{\rm NR}^{32}=\frac{\varepsilon_{1}-\varepsilon_{2}}{\varepsilon_{1}+\varepsilon_{2}}\times\frac{\varepsilon_{3}-\varepsilon_{2}}{\varepsilon_{3}+\varepsilon_{2}}\,.$ (4)
Hence, the Hamaker constant (in the non-retarded limit) can be expressed as,
$A^{\text{NR}}=\frac{3}{2}k_{B}T{\sum_{m=0}^{\infty}}{}^{\prime}\sum_{j=1}^{\infty}\Big[r_{\text{NR}}^{21}(i\xi_{m})r_{\text{NR}}^{23}(i\xi_{m})\Big]^{j}/j^{3}.$ (5)
There are crossovers between the product of the reflection coefficients in the non-retarded limit for different stoichiometry of the gapped Ca6-xAl7O16 metal surfaces in Fig. (3), that drives the possibility of attraction to repulsion transition in Casimir-Lifshitz interaction.
Figure 3: (Colors online) Product of the reflection coefficients at imaginary frequencies $\xi_{m}$ in non-retarded limit for ranges of gapped metals, Ca6-xAl7O16.
Hence, in a broad frequency range $\mathrm{r}_{\rm NR}^{12}\mathrm{r}_{\rm NR}^{32}>0$, the non-retarded Hamaker constant in Eq. (5) is positive, corresponding to an attractive interaction. However, other frequency intervals produce repulsive contributions.
Figure 4: (Colors online) The spectral functions $\mathrm{g}(\xi_{m})$, defined in Eq. (2), at different separations of liquid layer thickness in comparing the fully retarded and non-retarded limits in Ca6-xAl7O16-Methanol-PTFE like systems. Solid lines indicate the retardation effect, while dashed lines represent the behavior of spectral functions in the non-retarded limit.
In Fig. (4), we present the decomposition of the Casimir-Lifshitz free energy into distance and frequency-dependent spectral functions, $\mathrm{g}(\xi_{m})$. At large separations of liquid layer thickness, retardation effect influences the contributions from all frequencies except for the extremely low ones. Ultimately, as it is well known [1, 12, 32, 43], only the zero-frequency term survives. As the separation decreases, higher and higher frequencies play a role. As a result, the effect of retardation on repulsive contributions for the smallest separations is smaller since these originate from the low-frequency range. The key idea of this effect is a product of reflection coefficients (discussed in the previous part) along with a distance-dependent exponential factor $e^{-2\sqrt{{q}^{2}+\varepsilon_{2}\xi_{m}^{2}/c^{2}}d}$, meaning that significant contributions appear only when $\xi_{m}\sim\mathcal{O}(1/d)$ or smaller.
Figure 5: (Colors online) Free energy vs. distance plot in Ca6-xAl7O16 – Methanol – PTFE like system. Ca6-xAl7O16 indicates different stoichiometry. Here red color dot and magenta color box denote maximum value of Casimir energy for Ca6Al7O16 and Ca5.75Al7O16 are occurring at 2.9 nm and 9.49 nm respectively. Here solid lines represent interaction energy in full retarded limit where dashed lines point the interaction energy in non-retarded limit.
Figure 6: (Colors online) The retarded Casimir-Lifshitz pressure vs. distance plot in Ca6-xAl7O16 – Methanol – PTFE like system. Ca6-xAl7O16 indicates different stoichiometry. Negative pressure corresponds to attraction.
In Fig. (5), we notice that the non-retarded Casimir-Lifshitz interaction for different stoichiometry of Ca6-xAl7O16 is attractive in all separations, but there are reversals of signs for the retarded interaction around 2 nm and 5.9 nm for the metallic phases of the gapped metals Ca6Al7O16 and Ca5.75Al7O16, respectively. The sign reversal for pressure is shown in Fig. 6 where a positive pressure corresponds to repulsion. This sign reversal can also be viewed from zero-frequency Hamaker constant in Table (2). In order to demonstrate the long-range characteristics of various phases of gapped metals, the energy variation in the limit of long distances is illustrated in Fig. (7). It is observed that the metallic phases of gapped metals exhibit a repulsive effect even at long distances while the insulating phase displays an attractive interaction.
ine ine modes ($\omega_{j}$) Coefficients ($C_{j}$) for different Ca6-xAl7O16 compounds
ine Ca6Al7O16 Ca5.75Al7O16 Ca5.5Al7O16
ine 0.0206 58.9601 0.6494 0.0001
ine 0.0347 91.1774 1.797 0.0003
ine 0.0587 57.4068 5.2997 0.001
ine 0.1013 16.4729 15.4951 0.0221
ine 0.1996 73.0451 29.8463 0.3283
ine 0.3938 0.3949 2.1519 0.7511
ine 0.9556 0.0706 0.2519 0.4345
ine 2.2773 0.0987 0.0392 0.0
ine 6.4732 0.4594 0.2384 0.2279
ine 10.2048 0.7938 0.8189 0.0
ine 18.2421 0.3705 0.474 0.0486
ine 30.9018 0.1655 0.2388 0.0
ine 54.455 0.0059 0.0283 0.001
ine ine
[…]
ine Material $A^{NR}$ $A_{m=0}$
ine Ca6Al7O16 0.004 eV -0.012 eV
ine Ca5.75Al7O16 0.011 eV -0.004 eV
ine Ca5.5Al7O16 0.033 eV 0.016 eV
ine
Table 2: The non-retarded Hamaker constant for gapped metals–methanol–PTFE combinations.
Figure 7: (Colors online) Free energy vs.distance plot in Ca6-xAl7O16 – Methanol – Teflon like system in the long-distance limit. Ca6-xAl7O16 indicates different stoichiometry
The dielectric functions are larger for the metallic phases than for the insulating phase in the low-frequency range leading to the enhancement of repulsive contributions. In this figure, there is a smooth transition between the attractive van der Waals regime and the long-range Casimir region around 2 nm and 5.9 nm for the metallic phases of the gapped metals.
Remarkably, in most metallic cases the zero-frequency term dominates at 2.9 nm due to retardation and a strong cancellation of attractive and repulsive contributions. To the best of our knowledge, zero-frequency effects have never been predicted before, or observed, at such small distances at room temperature. For extremely short separations (less than $\sim$5-10 nm), the use of wave vector independent permittivities can still be a reasonable first order approximation. This was shown by Boström and Sernelius using the random phase approximation as a model for a q-dependent permittivity for thin and thick metallic films. [44]. Nonlocality is of course expected to influence the results in the limit of very short separations. [45, 46].
Figure 8: (Colors online) Decomposition of the contributions to the total non-retarded (non-ret) and retarded (ret) Casimir-Lifshitz free energy for Ca6Al7O16–Methanol–PTFE like system.
[…]
Figure 9: (Colors online) Free energy vs. distance plot in Ca6-xAl7O16 –Iodobenzene– PTFE like system. (a) In the inset, we illustrate the long-range characteristics of the Casimir-Lifshitz free energy, with an emphasis on the predominant contribution from zero-frequency term. The dotted black line represents zero free energy.
4 Conclusions
Here, we have predicted that how moving across regions with different off-stoichiometry can be used to tailor the sign of the Casimir-Lifshitz force between the Teflon nanosurface and gapped metal surfaces separated by methanol. Specifically, considering the range of gapped metals Ca6-xAl7O16, we demonstrate that controllable off-stoichiometry regions can be used to tune the Lifshitz interaction from attraction to repulsion. This behavior originates from the effect of off-stoichiometry on the dielectric function of gapped metals - the formation of cation vacancies induces a reduction of free carriers and the Drude contribution to the dielectric function. We show that the phase transition from insulator (Ca5.5Al7O16) to metal (Ca6Al7O16 and Ca5.75Al7O16) leads to a change in the force from attraction to repulsion at the nano to sub-micron scale driven by retardation and caused by the difference in dielectric functions of the gapped metals. The key point in our work is that a zero-frequency contribution and its interconnection with retardation and closeness of dielectric function of the liquids media and PTFE are the main origin for the sign change switching at unprecedented small distances. To the best of our knowledge, in the past purely zero-frequency effects have always been assumed to occur at much larger separations or at much higher temperatures [43]. Here, we see its domination for separations as small as a few nanometers in the metallic case but not for the insulating phase. While the dielectric functions used here do not have any temperature dependence (and hence can’t be used to represent the real temperature-dependent picture), one can find that, e.g., 10% changes in temperature, and the corresponding changes in the zero-frequency term, could lead to a measurable change in attraction-repulsion transition distances. With a change in background liquid media, it is also possible to have a zero-frequency Casimir-Lifshitz term leading to quantum levitation. In this way, using different liquids, we propose a new tool for nanoengineering via phase-transition-related quantum levitation in a liquid medium.
CRediT authorship contribution statement
Subhojit Pal: Writing – review -& editing, Writing – original draft, Software development, Investigation. Silvio Osella: Writing – review & editing. Oleksandr Malyi: Writing – review & editing, Materials modeling, Supervised computations, Computational resources acquisition. Mathias Boström: Writing – review -& editing, Writing – original draft, Initiated and supervised the project, Funding acquisition.
Declaration of Competing Interest
The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.
Acknowledgement
SP, OIM, and MB’s research contributions are part of the project No. 2022/47/P/ST3/01236 co-funded by the National Science Centre and the European Union’s Horizon 2020 research and innovation programme under the Marie Skłodowska-Curie grant agreement No. 945339. Institutional and infrastructural support for the ENSEMBLE3 Centre of Excellence was provided through the ENSEMBLE3 project (MAB/2020/14) delivered within the Foundation for Polish Science International Research Agenda Programme and cofinanced by the European Regional Development Fund and the Horizon 2020 Teaming for Excellence initiative (Grant Agreement No. 857543), as well as the Ministry of Education and Science initiative “Support for Centres of Excellence in Poland under Horizon 2020” (MEiN/2023/DIR/3797). S.O. thanks the National Science Centre, Poland (grant no. UMO/2020/39/I/ST4/01446) and the “Excellence Initiative - Research University” (IDUB) Program, Action I.3.3 - “Establishment of the Institute for Advanced Studies (IAS)” for funding (grant no. UW/IDUB/2020/25). We gratefully acknowledge Poland’s high-performance computing infrastructure PLGrid (HPC Centers: ACK Cyfronet AGH) for providing computer facilities and support within computational grant no. PLG/2023/016228 and for awarding this project access to the LUMI supercomputer, owned by the EuroHPC Joint Undertaking, hosted by CSC (Finland) and the LUMI consortium through grant no. PLL/2023/4/016319. M.B. gratefully acknowledges the kind hospitality at the University of Cagliari where this work was completed.